OpenStudy (anonymous):

The price of a shirt was increased from $20.00 to $30.00. What was the percent of increase?

OpenStudy (curry):

50%

OpenStudy (anonymous):

50%

OpenStudy (anonymous):

the increase is $10 and 10 is half of 20, so fifty percent

OpenStudy (anonymous):

20*1.5=30, so 50%

OpenStudy (gw2011):

The answer is 50% 30-20=10 10/20=0.50=50%

OpenStudy (anonymous):

the amount + %increase the amount = new amuont the amount=20 %increase = x new amount =30 so: 20 + 20x=30 20x=10 x=1/2=.5=50%

OpenStudy (anonymous):

sorry its %increase*amount

OpenStudy (anonymous):

30/20=1.5, the 0.5 is the percent increase 50%